OpenStudy (anonymous):

identify the coordinate pair of the point that partitions ab at a 1:5 ratio. round your answer to the nearest tenth. Coordinate A (8,3) coordinate B is (-2,-3)

OpenStudy (aum):



OpenStudy (anonymous):

I have the answers A. (7.3,2.6) B(9.7,4) C(6.3,2) D(-5.6,-2)

OpenStudy (aum):

AP:PB = 1:5 AP/AB = 1/6 Triangles APC and ABC are similar. Therefore, AP / AB = AC / AD = PC / BD = 1 / 6 AC / AD = 1 / 6 AC / -10 = 1 / 6 AC = -10 / 6 = -5/3 PC / BD = 1 / 6 PC / -6 = 1/ 6 PC = -1 X-coordinate of P = x-coordinate of A + AC = 8 - 5/3 = 19/3 = 6.3 Y-coordinate of P = y-coordinate of A + PC = 3 - 1 = 2 Point P is at (6.3, 2)

OpenStudy (anonymous):

thanks